OpenStudy (anonymous):

9-y^2 12y ------*----- 3-y 9+3y

OpenStudy (anonymous):

what do we hv to do here?

OpenStudy (anonymous):

you're multiplying the two fractions to equal something.

OpenStudy (anonymous):

then multiply the numerators and denominators separately and express as a fraction..thats it

OpenStudy (anonymous):

\[\frac{ 9-y ^{2} }{ 3-y } * \frac{ 12y }{ 9+3y }\]

OpenStudy (anonymous):

Yup!

OpenStudy (anonymous):

did you understand ineedromance? he is correct!

OpenStudy (anonymous):

\[\frac{ (9-y ^{2})(12y) }{ (3-y)(9+3y) }\]

OpenStudy (anonymous):

Thanks guys! I got it :)